OpenStudy (anonymous):

5. Which equation of a circle is represented by the graph? (Points : 1) x2 + y2 = 16 (x 1)2 + (y 1)2 = 4 (x 1)2 + (y 1)2 = 2 (x 1)2 + (y 1)2 = 16 i will atach doc

OpenStudy (anonymous):

here is the attachment

OpenStudy (anonymous):

D

OpenStudy (anonymous):

First, you have to understand the equation of a circle

OpenStudy (anonymous):

okay I hought it as the first one because the other ones apeard as lines when entered into my calculalotor

OpenStudy (anonymous):

could you help me with a few more?

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

thank you 1. Which equation of a circle is represented by the graph? (Points : 1) x2 + y2 = 16 (x 1)2 + (y 1)2 = 8 x2 + y2 = 8 (x 1)2 + (y 1)2 = 16

OpenStudy (anonymous):

i think it is the first one

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

3. Which of the graphs represents the circle with equation (x + 1)2 + (y 3)2 = 4? (Points : 1)

OpenStudy (anonymous):

Is it (y+3)^2 or (y-3)^2?

OpenStudy (anonymous):

the second y-

OpenStudy (anonymous):

and the 2's are ^2

OpenStudy (anonymous):

this one i think it is the last one but im not o sure if i am right.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

o cool, this is the last one _ Which is the graph of the circle with equation x2 + y2 = 9? (Points : 1)

OpenStudy (anonymous):

i dont know this one at all

OpenStudy (anonymous):

2nd one, I'll teach you how to do these

OpenStudy (anonymous):

yes please do lol im lost

OpenStudy (anonymous):

x^2+y^2, means the circle's center is at 0,0. Then, the square root of 9 tells you how long the radius is

OpenStudy (anonymous):

Do you have any other examples?

OpenStudy (anonymous):

These might help me explain to you

OpenStudy (anonymous):

not of the circles but I do of ones like these- 1. The circle with equation x2 + y2 = 100 is translated 2 units right and 3 units up. Where is the center of the new circle after the translation? (Points : 1) (2, 3) (3, 2) (2, 3) (3, 2)

OpenStudy (anonymous):

is that good? or do you need a circle one? if you do you can just use the last one we did i guess.

OpenStudy (anonymous):

Yeah, could you give the circle one please? That will help me explain better

OpenStudy (anonymous):

. Which equation of a circle is represented by the graph? (Points : 1) x2 + y2 = 16 (x 1)2 + (y 1)2 = 4 (x 1)2 + (y 1)2 = 2 (x 1)2 + (y 1)2 = 1

OpenStudy (anonymous):

Ok, first, identify the center point of the circle

OpenStudy (anonymous):

(1,1)?

OpenStudy (anonymous):

http://www.mathwarehouse.com/geometry/circle/equation-of-a-circle.php This site will explain better than i can lol

OpenStudy (anonymous):

cool, thank you for this site :)

OpenStudy (anonymous):

Your welcome :), it is an easy topic once you start to grasp the general idea of it, it is easier than graphing linear equations!!!

OpenStudy (anonymous):

If there is anything you do not understand, I will try to help you

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

Your welcome :)